# MnLargeSymbols’164 MnLargeSymbols’171

### v1

@@ 3.1 Lie dialgebra framework
$\widetilde{\mathfrak{g}}_{D}^{(0)}=\widetilde{\mathfrak{g}}_{0}^{\Gamma,0}\oplus\bigoplus_{r=1}^{N}\widetilde{\mathfrak{g}}_{r}\oplus\widetilde{\mathfrak{g}}_{\infty}^{\Gamma,0}\quad\text{and}\quad\widetilde{\mathfrak{g}}_{D}^{(1)}=\widetilde{\mathfrak{g}}_{0}^{\Gamma,1}\oplus\bigoplus_{r=1}^{N}\widetilde{\mathfrak{g}}_{r}\oplus\widetilde{\mathfrak{g}}_{\infty}^{\Gamma,1}\,,$ (3.13)
@@
$\begin{split}&\widetilde{\mathfrak{g}}_{0}^{\Gamma,k}=\{X_{0}(\lambda_{0})\in\widetilde{\mathfrak{g}}_{0}:\sigma(X_{0}(\lambda_{0}))=\omega^{k}X_{0}(\omega\lambda_{0})\}\,,\\ &\widetilde{\mathfrak{g}}_{\infty}^{\Gamma,k}=\{X_{\infty}(\lambda_{\infty})\in\widetilde{\mathfrak{g}}_{\infty}:\sigma(X_{\infty}(\lambda_{\infty}))=\omega^{k}X_{\infty}(\omega^{-1}\lambda_{\infty})\}\,,\end{split}$ (3.14)
@@
$X=\left(\sum_{n=-\infty}^{M_{0}-1}\frac{X_{0}^{(n)}}{\lambda^{n+1}},\sum_{n=-\infty}^{M_{1}-1}\frac{X_{1}^{(n)}}{(\lambda-\zeta_{1})^{n+1}},\ldots,\sum_{n=-\infty}^{M_{N}-1}\frac{X_{N}^{(n)}}{(\lambda-\zeta_{N})^{n+1}},\sum_{n=-\infty}^{M_{\infty}}\frac{X_{\infty}^{(n)}}{\lambda_{\infty}^{n}}\right)\,,$ (3.15)
@@
$Y=\left(\sum_{n=-\infty}^{N_{0}-1}\frac{Y_{0}^{(n)}}{\lambda^{n+1}},\sum_{n=-\infty}^{N_{1}-1}\frac{Y_{1}^{(n)}}{(\lambda-\zeta_{1})^{n+1}},\ldots,\sum_{n=-\infty}^{N_{N}-1}\frac{Y_{N}^{(n)}}{(\lambda-\zeta_{N})^{n+1}},\sum_{n=-\infty}^{N_{\infty}}\frac{Y_{\infty}^{(n)}}{\lambda_{\infty}^{n}}\right)\,,$ (3.16)
@@
The subspace $\widetilde{\mathfrak{g}}_{D}^{(0)}$ is a Lie subalgebra of $\widetilde{\mathfrak{g}}_{D}$ which defines the desired infinite-dimensional Lie algebra $\mathbb{g}=\widetilde{\mathfrak{g}}_{D}^{(0)}$ in the notation of Section 2. Continuing the identification of the different ingredients from the Lie dialgebra framework listed in Section 2, notice that $V=\widetilde{\mathfrak{g}}_{D}^{(1)}$ is clearly a representation of the Lie algebra $\mathbb{g}=\widetilde{\mathfrak{g}}_{D}^{(0)}$ since we have $[\widetilde{\mathfrak{g}}_{r}^{\Gamma,0},\widetilde{\mathfrak{g}}_{r}^{\Gamma,1}]\subset\widetilde{\mathfrak{g}}_{r}^{\Gamma,1}$ for $r\in\{0,\infty\}$. In the notation of Section 2, the representation $\operatorname{ad}^{\ast}:\mathbb{g}\times V\to V$ is given explicitly by $(X,Y)\mapsto[X,Y]$. The next proposition identifies $V=\widetilde{\mathfrak{g}}_{D}^{(1)}$ as a suitable model for the dual space $\mathbb{g}^{\ast}$.
@@
The bilinear pairing $\langle\cdot,\cdot\rangle:\widetilde{\mathfrak{g}}_{D}^{(1)}\times\widetilde{\mathfrak{g}}_{D}^{(0)}\to\mathbb{C}$ defined by
$\langle Y,X\rangle=T\sum_{r=1}^{N}\operatorname{Res}_{\lambda_{r}=0}\operatorname{Tr}(Y_{r}(\lambda_{r})X_{r}(\lambda_{r}))d\lambda+\sum_{r\in\{0,\infty\}}\operatorname{Res}_{\lambda_{r}=0}\operatorname{Tr}(Y_{r}(\lambda_{r})X_{r}(\lambda_{r}))d\lambda\,,$ (3.17)
@@
Next, we turn to the identification of the subalgebras $\mathbb{g}_{\pm}\subset\mathbb{g}$ and the corresponding subspaces $V_{\pm}\subset V$ in the notation of Section 2. Crucially for us, the spaces $\mathcal{F}_{D}^{\Gamma}$ and $\Omega_{D}^{\Gamma}$ embed into $\widetilde{\mathfrak{g}}_{D}^{(0)}$ and $\widetilde{\mathfrak{g}}_{D}^{(1)}$ respectively,
$\begin{split}&\iota_{\lambda}\colon\mathcal{F}_{D}^{\Gamma}\lhook\joinrel\relbar\joinrel\rightarrow\widetilde{\mathfrak{g}}_{D}^{(0)}\,,\quad f\longmapsto(\iota_{\lambda_{0}}f,\iota_{\lambda_{1}}f,\ldots,\iota_{\lambda_{N}}f,\iota_{\lambda_{\infty}}f)\,,\\ &\iota_{\lambda}\colon\Omega_{D}^{\Gamma}\lhook\joinrel\relbar\joinrel\rightarrow\widetilde{\mathfrak{g}}_{D}^{(1)}\,,\quad g\longmapsto(\iota_{\lambda_{0}}g,\iota_{\lambda_{1}}g,\ldots,\iota_{\lambda_{N}}g,\iota_{\lambda_{\infty}}g)\,,\end{split}$ (3.18)
@@
$\begin{split}&\pi_{\lambda}^{(0)}\colon\widetilde{\mathfrak{g}}_{D}^{(0)}\longrightarrow\mathcal{F}_{D}^{\Gamma}\,,\\ &\pi_{\lambda}^{(1)}\colon\widetilde{\mathfrak{g}}_{D}^{(1)}\longrightarrow\Omega_{D}^{\Gamma}\,,\end{split}$ (3.19)
@@
$\pi_{\lambda}^{(0)}\left(\sum_{n=-\infty}^{M_{0}-1}\frac{X_{0}^{(n)}}{\lambda^{n+1}},\sum_{n=-\infty}^{M_{1}-1}\frac{X_{1}^{(n)}}{(\lambda-\zeta_{1})^{n+1}},\ldots,\sum_{n=-\infty}^{M_{N}-1}\frac{X_{N}^{(n)}}{(\lambda-\zeta_{N})^{n+1}},\sum_{n=-\infty}^{M_{\infty}}\frac{X_{\infty}^{(n)}}{\lambda_{\infty}^{n}}\right)$
$\qquad\qquad\qquad\qquad\qquad=\sum_{n=0}^{M_{0}-1}\frac{X_{0}^{(n)}}{\lambda^{n+1}}+\sum_{r=1}^{N}\sum_{k=0}^{T-1}\sum_{n=0}^{M_{r}-1}\frac{\sigma^{k}X_{r}^{(n)}}{(\omega^{-k}\lambda-\zeta_{r})^{n+1}}+\sum_{n=0}^{M_{\infty}}\frac{X_{\infty}^{(n)}}{\lambda_{\infty}^{n}}\,,$ (3.20)
@@
$\pi_{\lambda}^{(1)}\left(\sum_{n=-\infty}^{N_{0}-1}\frac{Y_{0}^{(n)}}{\lambda^{n+1}},\sum_{n=-\infty}^{N_{1}-1}\frac{Y_{1}^{(n)}}{(\lambda-\zeta_{1})^{n+1}},\ldots,\sum_{n=-\infty}^{N_{N}-1}\frac{Y_{N}^{(n)}}{(\lambda-\zeta_{N})^{n+1}},\sum_{n=-\infty}^{N_{\infty}}\frac{Y_{\infty}^{(n)}}{\lambda_{\infty}^{n}}\right)$
$\qquad\qquad\qquad\qquad\qquad=\sum_{n=0}^{N_{0}-1}\frac{Y_{0}^{(n)}}{\lambda^{n+1}}+\sum_{r=1}^{N}\sum_{k=0}^{T-1}\sum_{n=0}^{N_{r}-1}\frac{\omega^{-k}\sigma^{k}Y_{r}^{(n)}}{(\omega^{-k}\lambda-\zeta_{r})^{n+1}}+\sum_{n=0}^{N_{\infty}}\frac{Y_{\infty}^{(n)}}{\lambda_{\infty}^{n}}\,.$ (3.21)
We will also need to define the following subspaces of $\widetilde{\mathfrak{g}}_{D}^{(0)}$ and $\widetilde{\mathfrak{g}}_{D}^{(1)}$ respectively:
$\widetilde{\mathfrak{g}}_{D+}^{(0)}=\widetilde{\mathfrak{g}}_{0+}^{\Gamma,0}\oplus\bigoplus_{r=1}^{N}\widetilde{\mathfrak{g}}_{r+}\oplus\widetilde{\mathfrak{g}}_{\infty+}^{\Gamma,1}\quad\text{and}\quad\widetilde{\mathfrak{g}}_{D+}^{(1)}=\widetilde{\mathfrak{g}}_{0+}^{\Gamma,1}\oplus\bigoplus_{r=1}^{N}\widetilde{\mathfrak{g}}_{r+}\oplus\widetilde{\mathfrak{g}}_{\infty+}^{\Gamma,1}$ (3.22)
@@
$\begin{split}&\widetilde{\mathfrak{g}}_{0+}^{\Gamma,k}=\widetilde{\mathfrak{g}}_{0}^{\Gamma,k}\cap\mathfrak{g}\otimes\mathbb{C}[[\lambda_{0}]]\,,\\ &\widetilde{\mathfrak{g}}_{r+}=\mathfrak{g}\otimes\mathbb{C}[[\lambda_{r}]]\,,\quad r\in\{1,\ldots,N\}\,,\\ &\widetilde{\mathfrak{g}}_{\infty+}^{\Gamma,k}=\widetilde{\mathfrak{g}}_{\infty}^{\Gamma,k}\cap\mathfrak{g}\otimes\lambda_{\infty}\mathbb{C}[[\lambda_{\infty}]]\,.\end{split}$ (3.23)
@@
Coming back to the identification of the Lie dialgebra ingredients from Section 2, the next proposition identifies the desired decomposition (2.3) of $\mathbb{g}=\widetilde{\mathfrak{g}}_{D}^{(0)}$ into complementary subalgebras $\mathbb{g}_{\pm}\subset\mathbb{g}$. Explicitly, we have the identifications $\mathbb{g}_{+}=\widetilde{\mathfrak{g}}_{D+}^{(0)}$ and $\mathbb{g}_{-}=\iota_{\lambda}\mathcal{F}_{D}^{\Gamma}$. We also identify a decomposition (2.4) of our model $V=\widetilde{\mathfrak{g}}_{D}^{(1)}$ for the dual space $\mathbb{g}^{\ast}$ into complementary subspaces $V_{\pm}\subset V$, where explicitly $V_{+}=\widetilde{\mathfrak{g}}_{D+}^{(1)}$ and $V_{-}=\iota_{\lambda}\Omega_{D}^{\Gamma}$, but will show only later in Proposition 3.4 that this decomposition of $V$ is the desired one induced by that of $\mathbb{g}$.
@@
$\widetilde{\mathfrak{g}}_{D}^{(0)}=\widetilde{\mathfrak{g}}_{D+}^{(0)}\dotplus\iota_{\lambda}\mathcal{F}_{D}^{\Gamma}\,,$ (3.24)
$\widetilde{\mathfrak{g}}_{D}^{(1)}=\widetilde{\mathfrak{g}}_{D+}^{(1)}\dotplus\iota_{\lambda}\Omega_{D}^{\Gamma}\,,$ (3.25)
respectively. Moreover, the subspaces $\widetilde{\mathfrak{g}}_{D+}^{(0)}$ and $\iota_{\lambda}\mathcal{F}_{D}^{\Gamma}$ are Lie subalgebras of $\widetilde{\mathfrak{g}}_{D}^{(0)}$.
@@
To any $X=(X_{0},X_{1},\ldots,X_{N},X_{\infty})\in\widetilde{\mathfrak{g}}_{D}^{(0)}$, we associate an equivariant function $f=\pi_{\lambda}^{(0)}X$. Notice that for all $r\in S$, $X_{r}-\iota_{\lambda_{r}}f$ is a formal Taylor series in $\lambda_{r}$. It follows that $X$ splits uniquely as the direct sum of the tuples $\left(X_{0}-\iota_{\lambda_{0}}f,X_{1}-\iota_{\lambda_{1}}f,\ldots,X_{N}-\iota_{\lambda_{N}}f,\right.$ $\left.X_{\infty}-\iota_{\lambda_{\infty}}f\right)\in\widetilde{\mathfrak{g}}_{D+}^{(0)}$ and $\left(\iota_{\lambda_{0}}f,\iota_{\lambda_{1}}f,\ldots,\iota_{\lambda_{N}}f,\iota_{\lambda_{\infty}}f\right)\in\iota_{\lambda}\mathcal{F}_{D}^{\Gamma}$, as required.
@@
Let $P_{+}$ and $P_{-}$ denote the projectors onto the subspaces $\widetilde{\mathfrak{g}}_{D+}^{(0)}$ and $\iota_{\lambda}\mathcal{F}_{D}^{\Gamma}$ respectively, relative to the decomposition in (3.24), and $\mathcal{P}_{+}$ and $\mathcal{P}_{-}$ the projectors onto $\widetilde{\mathfrak{g}}_{D+}^{(1)}$ and $\iota_{\lambda}\Omega_{D}^{\Gamma}$ respectively, relative to the decomposition in (3.25), in line with the notation from Section 2. As recalled in Section 2, the linear map
$R=P_{+}-P_{-}\,,$ (3.26)

### v2

@@ 3.1 Lie dialgebra framework
$\widetilde{\mathfrak{g}}_{D}^{(0)}=\widetilde{\mathfrak{g}}_{0}^{\Gamma,0}\oplus\bigoplus_{r=1}^{N}\widetilde{\mathfrak{g}}_{r}\oplus\widetilde{\mathfrak{g}}_{\infty}^{\Gamma,0}\qquad\text{and}\qquad\widetilde{\mathfrak{g}}_{D}^{(1)}=\widetilde{\mathfrak{g}}_{0}^{\Gamma,1}\oplus\bigoplus_{r=1}^{N}\widetilde{\mathfrak{g}}_{r}\oplus\widetilde{\mathfrak{g}}_{\infty}^{\Gamma,1},$
@@
$\widetilde{\mathfrak{g}}_{0}^{\Gamma,k}=\bigl{\{}X_{0}(\lambda_{0})\in\widetilde{\mathfrak{g}}_{0}\mid\sigma(X_{0}(\lambda_{0}))=\omega^{k}X_{0}(\omega\lambda_{0})\bigr{\}},$
$\widetilde{\mathfrak{g}}_{\infty}^{\Gamma,k}=\bigl{\{}X_{\infty}(\lambda_{\infty})\in\widetilde{\mathfrak{g}}_{\infty}\mid\sigma(X_{\infty}(\lambda_{\infty}))=\omega^{k}X_{\infty}\bigl{(}\omega^{-1}\lambda_{\infty}\bigr{)}\bigr{\}}$ (3.7)
@@
$X=\left(\sum_{n=-\infty}^{M_{0}-1}\frac{X_{0}^{(n)}}{\lambda^{n+1}},\sum_{n=-\infty}^{M_{1}-1}\frac{X_{1}^{(n)}}{(\lambda-\zeta_{1})^{n+1}},\dots,\sum_{n=-\infty}^{M_{N}-1}\frac{X_{N}^{(n)}}{(\lambda-\zeta_{N})^{n+1}},\sum_{n=-\infty}^{M_{\infty}}\frac{X_{\infty}^{(n)}}{\lambda_{\infty}^{n}}\right),$ (3.8)
@@
$Y=\left(\sum_{n=-\infty}^{N_{0}-1}\frac{Y_{0}^{(n)}}{\lambda^{n+1}},\sum_{n=-\infty}^{N_{1}-1}\frac{Y_{1}^{(n)}}{(\lambda-\zeta_{1})^{n+1}},\dots,\sum_{n=-\infty}^{N_{N}-1}\frac{Y_{N}^{(n)}}{(\lambda-\zeta_{N})^{n+1}},\sum_{n=-\infty}^{N_{\infty}}\frac{Y_{\infty}^{(n)}}{\lambda_{\infty}^{n}}\right),$ (3.9)
@@
The subspace $\widetilde{\mathfrak{g}}_{D}^{(0)}$ is a Lie subalgebra of $\widetilde{\mathfrak{g}}_{D}$ which defines the desired infinite-dimensional Lie algebra $\mathbb{g}=\widetilde{\mathfrak{g}}_{D}^{(0)}$ in the notation of Section 2. Continuing the identification of the different ingredients from the Lie dialgebra framework listed in Section 2, notice that $V=\widetilde{\mathfrak{g}}_{D}^{(1)}$ is clearly a representation of the Lie algebra $\mathbb{g}=\widetilde{\mathfrak{g}}_{D}^{(0)}$ since we have $\bigl{[}\widetilde{\mathfrak{g}}_{r}^{\Gamma,0},\widetilde{\mathfrak{g}}_{r}^{\Gamma,1}\bigr{]}\subset\widetilde{\mathfrak{g}}_{r}^{\Gamma,1}$ for $r\in\{0,\infty\}$. In the notation of Section 2, the representation $\operatorname{ad}^{\ast}\colon\mathbb{g}\times V\to V$ is given explicitly by $(X,Y)\mapsto[X,Y]$. The next proposition identifies $V=\widetilde{\mathfrak{g}}_{D}^{(1)}$ as a suitable model for the dual space $\mathbb{g}^{\ast}$.
@@
The bilinear pairing $\langle\cdot,\cdot\rangle\colon\widetilde{\mathfrak{g}}_{D}^{(1)}\times\widetilde{\mathfrak{g}}_{D}^{(0)}\to\mathbb{C}$ defined by
$\langle Y,X\rangle=T\sum_{r=1}^{N}\operatorname{Res}_{\lambda_{r}=0}\operatorname{Tr}(Y_{r}(\lambda_{r})X_{r}(\lambda_{r})){\rm d}\lambda+\sum_{r\in\{0,\infty\}}\operatorname{Res}_{\lambda_{r}=0}\operatorname{Tr}(Y_{r}(\lambda_{r})X_{r}(\lambda_{r})){\rm d}\lambda$ (3.10)
@@
Recall that the trace $\operatorname{Tr}\colon\mathfrak{g}\times\mathfrak{g}\to\mathbb{C}$, $(x,y)\mapsto\operatorname{Tr}(xy)$ is a nondegenerate invariant bilinear pairing on $\mathfrak{g}$ and invariant under the action of $\sigma$. Given any $x\in\mathfrak{g}^{(m)}$ and $y\in\mathfrak{g}^{(n)}$, we have $\operatorname{Tr}(xy)=\operatorname{Tr}(\sigma(x)\sigma(y))=\omega^{m+n}\operatorname{Tr}(xy)$. Therefore, $\operatorname{Tr}(xy)=0$ for all $x\in\mathfrak{g}^{(m)}$ and $y\in\mathfrak{g}^{(n)}$ if $m+n\neq 0\mod T$. Now, since $\operatorname{Tr}\colon\mathfrak{g}\times\mathfrak{g}\to\mathbb{C}$ is nondegenerate on $\mathfrak{g}$, given $y\in\mathfrak{g}^{(n)}$, there is $x\in\mathfrak{g}$ such that $\operatorname{Tr}(xy)\neq 0$. From (3.2), we have $x=x^{(0)}+\dots+x^{(T-1)}$, and from the previous result $\operatorname{Tr}(xy)=\operatorname{Tr}\bigl{(}x^{(m)}y\bigr{)}$ with $m+n=0\mod T$. Hence, $\operatorname{Tr}\bigl{(}x^{(m)}y\bigr{)}\neq 0$ and we get a nondegenerate pairing between the subspaces $\mathfrak{g}^{(m)}$ and $\mathfrak{g}^{(n)}$ with $m+n=0\mod T$.
@@
Next, we turn to the identification of the subalgebras $\mathbb{g}_{\pm}\subset\mathbb{g}$ and the corresponding subspaces $V_{\pm}\subset V$ in the notation of Section 2. Crucially for us, the spaces $\mathcal{F}_{D^{\prime}}^{\Gamma}$ and $\Omega_{D^{\prime}}^{\Gamma}$ embed into $\widetilde{\mathfrak{g}}_{D}^{(0)}$ and $\widetilde{\mathfrak{g}}_{D}^{(1)}$, respectively,
$\iota_{\lambda}\colon\ \mathcal{F}_{D^{\prime}}^{\Gamma}\lhook\joinrel\relbar\joinrel\rightarrow\widetilde{\mathfrak{g}}_{D}^{(0)},\qquad f\longmapsto(\iota_{\lambda_{0}}f,\iota_{\lambda_{1}}f,\dots,\iota_{\lambda_{N}}f,\iota_{\lambda_{\infty}}f),$
$\iota_{\lambda}\colon\ \Omega_{D^{\prime}}^{\Gamma}\lhook\joinrel\relbar\joinrel\rightarrow\widetilde{\mathfrak{g}}_{D}^{(1)},\qquad g\longmapsto(\iota_{\lambda_{0}}g,\iota_{\lambda_{1}}g,\dots,\iota_{\lambda_{N}}g,\iota_{\lambda_{\infty}}g),$ (3.11)
@@
$\pi_{\lambda}^{(0)}\colon\ \widetilde{\mathfrak{g}}_{D}^{(0)}\longrightarrow\mathcal{F}_{D^{\prime}}^{\Gamma},$
$\pi_{\lambda}^{(1)}\colon\ \widetilde{\mathfrak{g}}_{D}^{(1)}\longrightarrow\Omega_{D^{\prime}}^{\Gamma},$ (3.12)
@@
$\pi_{\lambda}^{(0)}\Biggl{(}\sum_{n=-\infty}^{M_{0}-1}\frac{X_{0}^{(n)}}{\lambda^{n+1}},\sum_{n=-\infty}^{M_{1}-1}\frac{X_{1}^{(n)}}{(\lambda-\zeta_{1})^{n+1}},\dots,\sum_{n=-\infty}^{M_{N}-1}\frac{X_{N}^{(n)}}{(\lambda-\zeta_{N})^{n+1}},\sum_{n=-\infty}^{M_{\infty}}\frac{X_{\infty}^{(n)}}{\lambda_{\infty}^{n}}\Biggr{)}$
$\qquad{}=\sum_{n=0}^{M_{0}-1}\frac{X_{0}^{(n)}}{\lambda^{n+1}}+\sum_{r=1}^{N}\sum_{k=0}^{T-1}\sum_{n=0}^{M_{r}-1}\frac{\sigma^{k}X_{r}^{(n)}}{(\omega^{-k}\lambda-\zeta_{r})^{n+1}}+\sum_{n=0}^{M_{\infty}}\frac{X_{\infty}^{(n)}}{\lambda_{\infty}^{n}},$
@@
$\pi_{\lambda}^{(1)}\Biggl{(}\sum_{n=-\infty}^{N_{0}-1}\frac{Y_{0}^{(n)}}{\lambda^{n+1}},\sum_{n=-\infty}^{N_{1}-1}\frac{Y_{1}^{(n)}}{(\lambda-\zeta_{1})^{n+1}},\dots,\sum_{n=-\infty}^{N_{N}-1}\frac{Y_{N}^{(n)}}{(\lambda-\zeta_{N})^{n+1}},\sum_{n=-\infty}^{N_{\infty}}\frac{Y_{\infty}^{(n)}}{\lambda_{\infty}^{n}}\Biggr{)}$
$\qquad{}=\sum_{n=0}^{N_{0}-1}\frac{Y_{0}^{(n)}}{\lambda^{n+1}}+\sum_{r=1}^{N}\sum_{k=0}^{T-1}\sum_{n=0}^{N_{r}-1}\frac{\omega^{-k}\sigma^{k}Y_{r}^{(n)}}{(\omega^{-k}\lambda-\zeta_{r})^{n+1}}+\sum_{n=0}^{N_{\infty}}\frac{Y_{\infty}^{(n)}}{\lambda_{\infty}^{n}}.$
We will also need to define the following subspaces of $\widetilde{\mathfrak{g}}_{D}^{(0)}$ and $\widetilde{\mathfrak{g}}_{D}^{(1)}$, respectively:
$\widetilde{\mathfrak{g}}_{D+}^{(0)}=\widetilde{\mathfrak{g}}_{0+}^{\Gamma,0}\oplus\bigoplus_{r=1}^{N}\widetilde{\mathfrak{g}}_{r+}\oplus\widetilde{\mathfrak{g}}_{\infty+}^{\Gamma,1}\qquad\text{and}\qquad\widetilde{\mathfrak{g}}_{D+}^{(1)}=\widetilde{\mathfrak{g}}_{0+}^{\Gamma,1}\oplus\bigoplus_{r=1}^{N}\widetilde{\mathfrak{g}}_{r+}\oplus\widetilde{\mathfrak{g}}_{\infty+}^{\Gamma,1},$
@@
$\widetilde{\mathfrak{g}}_{0+}^{\Gamma,k}=\widetilde{\mathfrak{g}}_{0}^{\Gamma,k}\cap\mathfrak{g}\otimes\mathbb{C}[[\lambda_{0}]],$
$\widetilde{\mathfrak{g}}_{r+}=\mathfrak{g}\otimes\mathbb{C}[[\lambda_{r}]],\qquad r\in\{1,\dots,N\},$
$\widetilde{\mathfrak{g}}_{\infty+}^{\Gamma,k}=\widetilde{\mathfrak{g}}_{\infty}^{\Gamma,k}\cap\mathfrak{g}\otimes\lambda_{\infty}\mathbb{C}[[\lambda_{\infty}]].$
@@
Coming back to the identification of the Lie dialgebra ingredients from Section 2, the next proposition identifies the desired decomposition (2.3) of $\mathbb{g}=\widetilde{\mathfrak{g}}_{D}^{(0)}$ into complementary subalgebras ${\mathbb{g}_{\pm}\subset\mathbb{g}}$. Explicitly, we have the identifications $\mathbb{g}_{+}=\widetilde{\mathfrak{g}}_{D+}^{(0)}$ and $\mathbb{g}_{-}=\iota_{\lambda}\mathcal{F}_{D^{\prime}}^{\Gamma}$. We also identify a decomposition (2.4) of our model $V=\widetilde{\mathfrak{g}}_{D}^{(1)}$ for the dual space $\mathbb{g}^{\ast}$ into complementary subspaces $V_{\pm}\subset V$, where explicitly $V_{+}=\widetilde{\mathfrak{g}}_{D+}^{(1)}$ and $V_{-}=\iota_{\lambda}\Omega_{D^{\prime}}^{\Gamma}$, but will show only later in Proposition 3.4 that this decomposition of $V$ is the desired one induced by that of $\mathbb{g}$.
@@
$\widetilde{\mathfrak{g}}_{D}^{(0)}=\widetilde{\mathfrak{g}}_{D+}^{(0)}\dotplus\iota_{\lambda}\mathcal{F}_{D^{\prime}}^{\Gamma},$ (3.13)
$\widetilde{\mathfrak{g}}_{D}^{(1)}=\widetilde{\mathfrak{g}}_{D+}^{(1)}\dotplus\iota_{\lambda}\Omega_{D^{\prime}}^{\Gamma},$ (3.14)
respectively. Moreover, the subspaces $\widetilde{\mathfrak{g}}_{D+}^{(0)}$ and $\iota_{\lambda}\mathcal{F}_{D^{\prime}}^{\Gamma}$ are Lie subalgebras of $\widetilde{\mathfrak{g}}_{D}^{(0)}$.
@@
To any $X\!=\!(X_{0},X_{1},\dots,X_{N},X_{\infty})\!\in\!\widetilde{\mathfrak{g}}_{D}^{(0)}$, we associate an equivariant function ${f\!=\!\pi_{\lambda}^{(0)}X}$. Notice that for all $r\in S$, $X_{r}-\iota_{\lambda_{r}}f$ is a formal Taylor series in $\lambda_{r}$. It follows that $X$ splits uniquely as the direct sum of the tuples $(X_{0}-\iota_{\lambda_{0}}f,X_{1}-\iota_{\lambda_{1}}f,\dots,X_{N}-\iota_{\lambda_{N}}f,X_{\infty}-\iota_{\lambda_{\infty}}f)\in\widetilde{\mathfrak{g}}_{D+}^{(0)}$ and $(\iota_{\lambda_{0}}f,\iota_{\lambda_{1}}f,\dots,\iota_{\lambda_{N}}f,\iota_{\lambda_{\infty}}f)\in\iota_{\lambda}\mathcal{F}_{D^{\prime}}^{\Gamma}$, as required.
@@
Let $P_{+}$ and $P_{-}$ denote the projectors onto the subspaces $\widetilde{\mathfrak{g}}_{D+}^{(0)}$ and $\iota_{\lambda}\mathcal{F}_{D^{\prime}}^{\Gamma}$ respectively, relative to the decomposition in (3.13), and $\mathcal{P}_{+}$ and $\mathcal{P}_{-}$ the projectors onto $\widetilde{\mathfrak{g}}_{D+}^{(1)}$ and $\iota_{\lambda}\Omega_{D^{\prime}}^{\Gamma}$ respectively, relative to the decomposition in (3.14), in line with the notation from Section 2. As recalled in Section 2, the linear map
$R=P_{+}-P_{-}$ (3.15)